OpenStudy (anonymous):

The perimeter of a rectangle is 36 ft. The length is 10 ft longer than the width. find the dimensions. write a system of oinear equations and solve the resulting system. 2x+2y=

OpenStudy (anonymous):

i gave you an answer already down there.

OpenStudy (anonymous):

let w = width; L=length; P=2W+2L L=W+10 substitute this for L and 36 for P 36=2W+2(W+10) 36=2W+2W+20 [distribute] 16=4W [collect like terms; subtract 20] 4=W [divide by 4] L=4+10=14